OpenStudy (anonymous):

I will give medal and fan. HELP URGENT Nicole rented a boat at $325 for 2 days. If she rents the same boat for 5 days, she has to pay a total rent of $892. Part A: Write the information above in point slope form: Write this equation in slope intercept form: Part B: Write the equation in part A in standard form (2 points) Part C: Describe the steps to graph the equation in Part A. What is the label of the x-axis? What is the label of the y-axis? What is the slope? What is the y- intercept? What is the x-intercept?

OpenStudy (aum):

We can treat the cost as the y-value and the number of days as the x-value. We are given two points. when x = 2, y = $325; when x = 5, y = $892 Slope = \(\Large \frac{y_2 - y_1}{x_2-x_1} = \frac{892-325}{5-2} = \frac{567}{3} = 189\) Point Slope form: \(y - y_1 = m(x-x_1)\) \(y - 325 = 189(x-2)\)

OpenStudy (anonymous):

ok sorry its a lot of work for the problem

OpenStudy (aum):

y - 325 = 189(x - 2) y - 325 = 189x - 378 y = 189x - 378 + 325 y = 189x - 53 is the slope-intercept form.

OpenStudy (aum):

B) y = 189x - 53 53 + y = 189x 53 = 189x - y 189x - y = 53 is the standard form.

OpenStudy (aum):

Part C: Describe the steps to graph the equation in Part A. Cost is chosen as the y-value and the number of days is the x-value. We are given two points: (2, 325) and (5, 892). We choose the x-scale as 1 unit = 1 day. We choose the y-scale as 1 unit = $100. We plot the two points on the graph paper. We draw a straight line through the two points for positive x and positive y values. What is the label of the x-axis? Days What is the label of the y-axis? Cost in Dollars What is the slope? 189 What is the y- intercept? y = 189x - 53. Put x = 0: y = -53. So the y-intercept is -53. What is the x-intercept? y = 189x - 53. Put y = 0: 0 = 189x - 53. x = 53/189 So the y-intercept is 53/189.

OpenStudy (anonymous):

thank you so so so so much you're the best I would have been up till midnight trying to figure this out thank you

OpenStudy (aum):

You are welcome.

OpenStudy (aum):

The last line should be: The x-intercept is 53/189 (not y-intercept).